# Token-Aware Coding Flow: A Study with Nano Surge in Reasoning Model thanks: * Corresponding authors

### v1

@@ Abstract
With the widespread application of large-scale language models (LLMs) in software engineering, the Chain of Thought (CoT) approach has emerged as a crucial tool for driving automated code generation and optimization. However, despite the significant success of CoT methods in generating high-quality code, the issue of token inflation during the reasoning process remains a formidable challenge to model performance and efficiency, particularly when dealing with complex code smells. Code smells not only affect the maintainability and scalability of code but also significantly increase the computational burden during LLM inference, leading to excessive token consumption and, consequently, reduced reasoning efficiency. This paper introduces an innovative Token-Aware Coding Flow method, aimed at addressing the token inflation problem caused by smelly code in the CoT process. Through experimentation, we validate the synergistic effect of code refactoring and prompt engineering strategies, demonstrating that after eliminating code smells, token consumption during model inference is significantly reduced. The experimental results show that refactored code, while maintaining functional consistency, can reduce token consumption by up to 50%. Additionally, by explicitly prompting the type of code smells in the prompt and incorporating strategies such as context awareness and role constraints, we further optimize the reasoning process, achieving a 24.5% to 30% reduction in token consumption. These optimizations not only significantly enhance the model’s reasoning efficiency and improve code generation quality but also provide new insights for addressing performance bottlenecks in complex code generation tasks.
@@ I Introduction
Figure 1: Evolution of software development paradigms. (1) Traditional coding paradigm: human programmers manually translate requirements into code. (2) Augmented coding paradigm: programmers are assisted by base language models (e.g., GPT-3.5, GitHub Copilot) to generate code, reducing manual effort. (3) Reasoning-driven paradigm: advanced reasoning models (e.g., GPT-o3, DeepSeek-R1) interpret requirements and autonomously generate code by simulating human-like thinking, leading to deeper understanding and improved code quality.
With the rapid proliferation of large language models (LLMs), such as GPT-o3 and DeepSeek-R1[1], in the field of software development[2], traditional software development paradigms have undergone profound changes(Fig. 1). In the past, software functionality was implemented solely through manual programming; however, it is now gradually shifting toward human-machine collaboration[3], where code generation and optimization are assisted or directly achieved through the ”Chain of Thought” (CoT) approach[4]. This new development paradigm significantly improves both development efficiency and code quality[5]. However, existing research primarily focuses on LLM performance under ideal conditions, such as code that is well-structured and well-designed[6], neglecting the impact of ”code smells” that are commonly present in industrial settings due to legacy issues or insufficient experience.
Code smells[7] are pervasive in industrial software systems and typically manifest as complex, redundant logical structures, ambiguous naming conventions, excessively long functions or parameter lists, and overly coupled class structures. While these issues may not immediately cause software failures, they severely impair code readability and maintainability in the long run, and may even lead to latent defects[8]. Previous studies have shown that human developers face an additional cognitive burden when dealing with these smelly codes, leading to reduced development efficiency[9]. However, it remains unclear how such smelly code affects model inference efficiency and token consumption, especially in the emerging LLM-driven development paradigm.
Preliminary practical observations suggest that large language models exhibit noticeable ”token inflation”[10] and ”smell regurgitation”[11] phenomena when processing code snippets with smells(Fig. 2). Specifically, the model may repeatedly validate and reason about a particular code fragment or logical unit during inference, resulting in an overall redundant reasoning process that significantly increases token consumption[12]. This not only raises the economic cost of using large models [13] but also severely limits their applicability in resource-constrained environments (e.g., mobile platforms and real-time applications). Therefore, a deeper understanding of and strategies to reduce the negative impact of code smells on model inference efficiency is of great practical significance.
@@
To further understand and attempt to address the above challenges, this paper proposes a method called Token-Aware Coding Flow. We first constructed a dataset containing ten common Java code smells, as well as a clean code dataset, based on a portion of the CodeXGLUE[14] Text-Code dataset, using the tree_sitter syntax tree construction tool[15]. This dataset was used to observe the token consumption differences when the DeepSeek-R1 model is faced with ”clean code” versus ”smelly code.” The results show that ”smelly code” consumes significantly more tokens per unit time than ”clean code,” highlighting the negative impact of code smells on model inference efficiency[16].
Building on this, we designed a series of exploratory experiments to systematically investigate the impact of different optimization strategies on token consumption. This study focuses on the impact of smelly code on inference models and introduces the ”Token-Aware Coding Flow” method. Through a series of carefully designed empirical experiments, we delve into the token consumption characteristics of smelly code during the Chain of Thought process and propose corresponding optimization strategies[17]. Specifically, this study aims to answer the following research questions (RQs):
Through experiments, we obtained the following key findings:
  • Finding1: Compared to clean code, the model consumes approximately 32% more tokens per unit time when processing smelly code.
  • Finding2: After eliminating code smells through refactoring, and ensuring functional consistency, token consumption during inference can be reduced by approximately 50%.
  • Finding3: The impact of different types of code smells varies significantly, with design and structural smells imposing a greater burden on token consumption than naming-related smells.
  • Finding4: Explicitly indicating code smell types in the prompt can reduce token consumption by approximately 24.5%.
  • Finding5: Combining context awareness, role constraints, and cost-sensitive prompt strategies can further reduce token consumption by 2% to 25%.
Based on the above research questions and findings, the main contributions of this paper are as follows:
  • Contribution1: We are the first to explore the impact of code smells on token consumption during model inference.
  • Contribution2: We validate the potential of code refactoring in reducing token consumption and provide a reusable method for automating the refactoring of smelly code.
  • Contribution3: We reveal the differential impact of various types of code smells on the model’s inference process.
  • Contribution4: We propose and validate effective prompt engineering strategies for smelly code handling.
II Context
II-A Code Smell
”Code Smell” refers to code characteristics that, while functional, may indicate underlying issues such as poor quality, maintenance difficulties, and scalability problems. These smells, like duplicated code, long methods, and excessive parameters[18], often result from bad programming practices and can negatively affect performance and readability. While they do not always cause immediate errors, they can complicate future maintenance. Developers often address code smells through refactoring, though manual inspection in large projects can be time-consuming and error-prone. Automated methods using machine learning and pre-trained models are increasingly being explored to improve the detection and elimination of code smells.
II-B Reasoning Module
The reasoning module is essential in intelligent code analysis and generation systems. It uses logical reasoning to interpret input (like code or task descriptions) and generate appropriate outputs. In code generation and repair tasks, it helps the model understand code structure, identify issues, and generate conforming code. With advancements in deep learning, reasoning modules have evolved from rule-based systems to those powered by large pre-trained models like GPT-3 and Codex. These models enhance reasoning accuracy and efficiency, making them crucial for tasks like automated code repair and generation(Fig. 3)[19].
@@ III-A Research Questions
Through these questions, we aim to gain a deeper understanding of the impact of smelly code on large model inference and explore the potential improvements in reasoning efficiency with different optimization strategies[20]. First, RQ1 primarily focuses on whether smelly code causes large models to consume more tokens during inference, which is crucial for evaluating the model’s ability to handle non-standard code. RQ2 investigates whether refactoring smelly code improves the model’s reasoning efficiency, particularly in terms of reducing token consumption while maintaining functional consistency. RQ3 focuses on the impact of different types of code smells on model reasoning efficiency, helping us understand which types of code smells have a more significant effect on token consumption. RQ4 examines whether explicitly pointing out the type of code smells in the prompt provides the model with more precise contextual information, thereby reducing redundant reasoning steps and lowering token consumption. Finally, RQ5 explores whether other effective prompt engineering strategies, such as context awareness, role constraints, and cost-sensitive strategies, can further optimize reasoning efficiency beyond code refactoring. By addressing these questions, we aim to both validate the specific impact of code smells on large model reasoning performance and provide experimental evidence for improving inference efficiency through refactoring and optimized prompt design.
@@ III-B1 Observational Experiment
Datasets: We used the Java dataset (Text-To-Code) from CodeXGLUE[14] for the experimental data, which includes 300 samples of smelly code and 300 samples of clean code. These datasets were used to analyze the behavior of the DeepSeek-R1 reasoning model when processing these codes, particularly the differences in token consumption and time consumption.
Metrics: We recorded the total number of tokens consumed by the model for each code and the time taken for reasoning. The token consumption per unit time was calculated as the core metric to assess the efficiency of large models when handling different types of code. Additionally, the generated code was evaluated based on functionality correctness, readability, robustness, and other indicators.
III-B2 Refactoring[22] and Optimization Experiment
Datasets: In this experiment, we used the dataset based on smelly code and applied DeepSeek-R1 for code refactoring to generate refactored code (rf_code). The refactored code was used for chain processing to compare its reasoning efficiency with the original smelly code and clean code.
Metrics: We recorded the number of tokens consumed in each processing step and calculated the Halstead complexity and cyclomatic complexity. By assessing the token consumption per unit of Halstead complexity and per line of code, we evaluated the efficiency improvements of the refactored code during model inference. Further, the accuracy of the generated code was assessed through functionality consistency scores (e.g., CodeBLEU) to ensure that the refactored code maintains functional consistency with the original code.
@@ III-B3 Code Smell Type Comparison Experiment
Datasets: We selected code samples from 10 different types of code smells in the smelly code dataset, with 50 samples from each type, totaling 500 samples. These samples were used to analyze the impact of different types of code smells on token consumption during model inference.
Metrics: By comparing the token consumption during reasoning for different types of code smells, we calculated the token consumption per unit time for each smell type. The differences in the impacts of design, structural, and naming-related smells were further analyzed. All generated code was evaluated for functional consistency to ensure that the model-generated code met the expected functionality during reasoning.
@@ III-B4 Explicit Smell Type Prompt Experiment
Datasets: This experiment was conducted on the previously used smelly code dataset, primarily comparing the token consumption differences when explicitly indicating code smell types in the prompt versus not indicating them.
Metrics: We recorded the number of tokens consumed in each experimental step and calculated token consumption per unit time. After generating the code, functionality consistency was assessed using metrics like CodeBLEU. Additionally, the effect of explicitly indicating code smell types was evaluated based on the model’s reasoning time, token consumption, and output quality.
@@ III-B5 Prompt Optimization Strategy Experiment
Datasets: In this experiment, we tested different prompt optimization strategies on the smelly code dataset[23], analyzing the effects of Context Awareness, Responsibility Tuning, and Cost-Sensitive strategies.
Metrics: We recorded token consumption during the reasoning process for each strategy and calculated the Halstead complexity and cyclomatic complexity for each code. By comparing token consumption per unit complexity and per line of code, we assessed the impact of the three prompt optimization strategies on reasoning efficiency. All experiments were also evaluated for functionality consistency and code generation quality to ensure the generated code was accurate[16].
@@ III-C Implementation
  • Data Processing [20] and Smell Annotation: Java code was extracted from the Text‑Code subset of CodeXGLUE and parsed with Tree‑Sitter to detect and classify instances of code smell. Leveraging Tree‑Sitter’s high‑performance syntax analysis, we annotated ten common smell categories, thereby constructing parallel “smelly” and “clean” code datasets. For each smelly sample, we invoked the DeepSeek‑R1 API to execute the reasoning task, recording both token consumption and inference latency. We then applied DeepSeek‑R1’s refactoring API to eliminate smells—producing rf_code—and compared the inference efficiency (tokens consumed and time taken) between original and refactored versions.
  • Prompt Engineering and Optimization[25]:To further reduce token overhead, we evaluated three prompt‑level strategies: context awareness, role constraints, and cost sensitivity. Each strategy was realized by modifying the model’s input prompt, thereby influencing the reasoning trajectory and token usage. For every experimental condition, we issued the corresponding API calls to adjust generation parameters and logged token counts, inference durations, and the resulting code quality.
  • Automation, Logging, and Analysis:All API interactions were fully automated via Python scripts, ensuring consistent and efficient execution. Each experiment’s outputs—including raw logs of token usage, timing, and generated code—were aggregated and processed by custom analysis scripts. These scripts perform data cleansing, statistical aggregation, and visualization, enabling systematic comparison across differing experimental configurations.
IV Experimental Results
IV-1 RQ1: Smelly Code vs. Clean Code

### v2

# Optimizing Token Consumption in LLMs: A Nano Surge Approach for Code Reasoning Efficiency thanks: * Corresponding authors

@@ Abstract
With the increasing adoption of large language models (LLMs) in software engineering, the Chain of Thought (CoT) reasoning paradigm has become an essential approach for automated code repair. However, the explicit multi-step reasoning in CoT leads to substantial increases in token consumption, reducing inference efficiency and raising computational costs, especially for complex code repair tasks. Most prior research has focused on improving the correctness of code repair while largely overlooking the resource efficiency of the reasoning process itself. To address this challenge, this paper proposes three targeted optimization strategies: Context Awareness, Responsibility Tuning, and Cost Sensitive. Context Awareness guides the model to focus on key contextual information, Responsibility Tuning refines the structure of the reasoning process through clearer role and responsibility assignment, and Cost Sensitive incorporates resource-awareness to suppress unnecessary token generation during inference. Experiments across diverse code repair scenarios demonstrate that these methods can significantly reduce token consumption in CoT-based reasoning without compromising repair quality. This work provides novel insights and methodological guidance for enhancing the efficiency of LLM-driven code repair tasks in software engineering.
@@ I Introduction
The rapid proliferation of Large Language Models (LLMs) is driving a profound transformation within the software engineering domain. From code generation and documentation to intricate system design and defect analysis, LLMs exhibit unprecedented potential, fundamentally altering traditional software development practices. Among these advancements, Chain of Thought (CoT)[1] has garnered widespread attention as a strategy designed to enhance the reasoning capabilities of these models. By prompting LLMs to emulate a human-like, step-by-step problem-solving process, CoT significantly improves model accuracy and output quality when tackling complex software engineering tasks, such as generating intricate code, remediating deep-seated errors, and executing multi-step optimizations[2][3]. This progress signifies a crucial shift for LLMs, from mere ”pattern matchers” to more sophisticated ”intelligent reasoners,” thereby unlocking new vistas for automated and intelligent software development[4].
However, the enhanced capabilities afforded by CoT are accompanied by a significant and often overlooked challenge: a substantial increase in token consumption[5]. The CoT mechanism inherently relies on the generation of detailed intermediate reasoning steps, which implies that both the length of input prompts and the volume of model-generated content can far exceed those of traditional, more succinct question-answering paradigms[6]. This phenomenon, termed ”token inflation,”[7] has emerged as a critical bottleneck impeding the broader application of LLMs in the software engineering field. When LLMs are tasked with analyzing large-scale, complex modern software systems, or engaged in programming tasks requiring multiple iterative refinements, the token inflation precipitated by CoT becomes particularly acute[8]. This not only imposes economic pressure on the cost-effectiveness of model utilization but also adversely impacts practical application efficiency.
In real-world software development applications, cost control and operational efficiency remain paramount considerations[9]. For all software development teams—ranging from large technology corporations and agile small-to-medium-sized enterprises to budget-constrained startups—the adoption of new technologies transcends mere enhancement of technical capabilities, necessitating a thorough evaluation of cost-effectiveness[6]. Currently, a majority of LLM services are accessed via API calls, with charges typically levied based on the volume of tokens processed. While CoT can yield higher-quality outputs, the concomitant substantial increase in token consumption may render its use prohibitively expensive[8]. Within modern software development pipelines, particularly in contexts such as mobile application development, embedded systems, and other resource-constrained environments, excessive token consumption can lead to considerable economic burdens, compelling many teams to exercise caution or curtail their reliance on CoT methodologies[1].
Beyond direct economic costs, excessive token consumption engenders other detrimental effects, including prolonged model response times and heightened demands on computational resources. Such issues can impair the agility of software development processes and the immediacy of user experiences, especially in scenarios mandating rapid responses and efficient delivery[6]. Consequently, achieving an optimal balance between harnessing the reasoning advantages of CoT and ensuring the economic viability and operational efficiency of LLM applications has surfaced as a significant challenge confronting the software engineering field[10].
This study does not aim to propose a systemic solution; rather, it represents a preliminary exploration of this issue. Through the observation and analysis of token consumption phenomena within CoT processes, we aspire to furnish insights for future research, investigating methodologies to effectively curtail token consumption without a significant detriment to reasoning performance. This research endeavors to pave the way for the expanded application of LLMs in software engineering, particularly within contexts that prioritize developmental efficiency and resource optimization.
Driven by widespread concerns regarding the token consumption induced by Chain of Thought (CoT) and its potential ramifications, this study undertakes an exploratory investigation. Our objective is to meticulously observe and comprehend the token dynamics of CoT when processing code of varying complexities within specific experimental settings[11]. Furthermore, we aim to identify external intervention strategies capable of effectively mitigating such consumption. To ground our research in realistic and common software development contexts—scenarios where the intricacies of CoT’s reasoning pathways and token expenditure may be amplified—we concentrate our observational lens on how Large Language Models (LLMs) handle code exhibiting different intrinsic characteristics. Specifically, the presence and nature of ”code smells[12][13]” are employed as a tangible proxy and experimental variable for examining these phenomena. It is crucial to emphasize that our intent is not to present ”code smell elimination” as an end in itself, but rather to utilize it as an effective ”lens” through which to gain deeper insights into the broader challenge of CoT token optimization[14].
Guided by this philosophy, we propose a conceptual methodological framework termed ”Token-Aware Coding Flow.” Based on this framework, we have conducted a series of preliminary experimental inquiries. These inquiries endeavor to offer initial insights into several key questions:
  • What is the specific impact of intrinsic code characteristics (e.g., as studied through a comparative analysis of ”clean code” versus ”smelly code”) on token consumption during CoT reasoning processes?
  • To what extent does pre-processing code (for instance, by refactoring to eliminate code smells, thereby simplifying the input furnished to CoT) demonstrate potential for reducing token consumption?
  • Do different types of inherent code complexities (exemplified by various categories of code smells) lead to differentiated patterns in token consumption?
  • When interacting with LLMs, does explicitly indicating potential code characteristics (such as the type of code smells) within the prompt help guide the CoT process toward more economical token usage?
Our preliminary experiments reveal a clear correlation between the intrinsic characteristics of code and token consumption during Chain of Thought (CoT) reasoning. Compared to ”clean” code, code with ”smells” significantly increases token usage, while refactoring to remove such code smells can substantially reduce the tokens required for subsequent reasoning. Furthermore, different types of code complexity (such as various code smells) affect token consumption differently, with deeper logical or structural issues typically imposing a higher token burden. Explicitly indicating code characteristics in the prompt, combined with prompt engineering strategies, also demonstrates potential for optimizing token efficiency.
Although this research is still in an exploratory phase and its conclusions require further validation, our initial results provide new perspectives on understanding and optimizing token consumption in CoT scenarios.Our main contributions of this work are as follows:
  • For the first time, we systematically focus on resource (token) consumption in code generation and processing tasks, drawing attention to the economic and efficiency impact of CoT reasoning in LLM-based software engineering.
  • We propose a novel approach to reduce resource consumption from outside the model, by leveraging code refactoring and prompt engineering techniques. This strategy offers practical and effective ways to optimize token usage without modifying model internals.
  • We empirically validate that code smells lead to significantly higher token consumption, and demonstrate that external optimization strategies (such as code refactoring and prompt design) can effectively reduce this overhead.
  • Our work provides empirical evidence and methodological inspiration for developing more cost-effective and efficient LLM-assisted software engineering workflows, contributing to the advancement of resource-aware intelligent development practices.
II Introduce the New Coding Flow
Figure 2: The Ladder of Programming and AI. (1) Traditional coding paradigm: human programmers manually translate requirements into code. (2) Augmented coding paradigm: programmers are assisted by base language models (e.g., GPT-3.5, GitHub Copilot) to generate code, reducing manual effort. (3) Reasoning-driven paradigm: advanced reasoning models (e.g., GPT-o3, DeepSeek-R1) interpret requirements and autonomously generate code by simulating human-like thinking, leading to deeper understanding and improved code quality.
The history of software development is characterized by a continuous pursuit of enhanced efficiency, superior quality, and greater automation capabilities. To better contextualize the significance of our proposed ”Token-Aware Coding Flow,” we first review and analyze the evolution of software development paradigms, particularly highlighting the recent transformations driven by Large Language Models (LLMs). As depicted in Figure 2, the evolution of software development models can be broadly categorized into the following three core stages[15][16][17]:
II-A Manual Programming
In the early phase of software development (Stage 1, Fig. 2), the process relied almost entirely on manual efforts by programmers. Every step—from requirements to design, implementation, testing, and maintenance—depended on their expertise and time. While this approach maximized human creativity, it suffered from low efficiency, limited scalability, difficult knowledge transfer, and a high risk of human errors and inconsistencies.
II-B Assisted Programming
Advancements such as IDEs and code assistance tools ushered in an assisted programming era (Stage 2, Fig. 2), where programmers leveraged technologies like auto-completion, syntax highlighting, version control, and basic code generators to boost productivity. More recently, LLM-powered tools (e.g., GPT-3.5, GitHub Copilot) have enabled smarter code suggestions, test generation, and documentation drafts. As a result, the programmer’s role shifted towards supervision and validation, marking the start of true human–AI collaboration.
II-C Cognitive Programming
Today, development is increasingly driven by advanced reasoning models (Stage 3, Fig. 2). Powerful LLMs with Chain of Thought (CoT) capabilities (such as GPT-o1, DeepSeek-R1[1], GPT-4, DeepSeek Coder) can understand, decompose, and autonomously address complex requirements—planning, generating, and optimizing multi-step code, and even participating in architecture and debugging. Developers now focus on defining goals and making decisions, while models handle much of the cognitive and implementation workload. This paradigm greatly enhances efficiency and problem-solving, potentially transforming software engineering.
However, as discussed in the Introduction, this reasoning-driven paradigm—especially its reliance on CoT—also leads to a sharp rise in token consumption. Optimizing token efficiency is thus essential to fully realize its benefits and ensure sustainable use in diverse development scenarios. This challenge motivates our proposal of the “Token-Aware Coding Flow” and the exploratory work in this study.
@@ III-A Research Questions
Through these questions, we aim to gain a deeper understanding of the impact of smelly code on large model inference and explore the potential improvements in reasoning efficiency with different optimization strategies[18]. First, RQ1 primarily focuses on whether smelly code causes large models to consume more tokens during inference, which is crucial for evaluating the model’s ability to handle non-standard code. RQ2 investigates whether refactoring smelly code improves the model’s reasoning efficiency, particularly in terms of reducing token consumption while maintaining functional consistency[6][19]. RQ3 focuses on the impact of different types of code smells on model reasoning efficiency, helping us understand which types of code smells have a more significant effect on token consumption[19]. RQ4 examines whether explicitly pointing out the type of code smells in the prompt provides the model with more precise contextual information, thereby reducing redundant reasoning steps and lowering token consumption[19]. Finally, RQ5 explores whether other effective prompt engineering strategies, such as context awareness, role constraints, and cost-sensitive strategies, can further optimize reasoning efficiency beyond code refactoring. By addressing these questions, we aim to both validate the specific impact of code smells on large model reasoning performance and provide experimental evidence for improving inference efficiency through refactoring and optimized prompt design.
@@ III-B1 Observational Experiment
Datasets: We used the CodeXGLUE Java (Text-To-Code) dataset, containing 300 smelly code samples and 300 clean code samples, to analyze differences in token and time consumption when DeepSeek-R1 processes different types of code.
Metrics: We recorded the total number of tokens and inference time for each code sample. Token consumption rate was calculated to evaluate the model’s efficiency in handling different code types. The generated code was further assessed for functional correctness and readability.
III-B2 Refactoring and Optimization Experiment
Datasets: Smelly code samples were refactored using DeepSeek-R1 to generate refactored code, which was then compared with the original smelly code and clean code to assess inference efficiency.
Metrics: We tracked token consumption at each processing step and calculated Halstead and cyclomatic complexity. Token consumption per unit of complexity and per line of code was analyzed. Functional consistency was ensured using metrics such as CodeBLEU.
@@ III-B3 Code Smell Type Comparison Experiment
Datasets: We selected 10 types of code smells from the smelly code dataset, with 50 samples per type (500 samples in total), to analyze the impact of different smell types on token consumption.
Metrics: Token consumption during inference was compared across different smell types, with additional analysis of design, structural, and naming smells. The generated code was also evaluated for functional consistency.
@@ III-B4 Explicit Smell Type Prompt Experiment
Datasets: Based on the smelly code dataset, we compared token consumption when the prompt explicitly indicated the code smell type versus when it did not.
Metrics: Token and time consumption were recorded at each experimental step. Functional consistency of the output was evaluated using metrics such as CodeBLEU. The impact of explicitly indicating the code smell type on inference efficiency and code quality was also analyzed.
@@ III-B5 Prompt Optimization Strategy Experiment
Datasets: Various prompt optimization strategies—such as context enhancement, responsibility tuning, and cost-sensitive prompting—were tested on the smelly code dataset.
Metrics: Token consumption and code complexity under each strategy were recorded. Token consumption per unit complexity and per line of code was compared. Each strategy was evaluated based on functional consistency and code generation quality.
@@ III-C Implementation
  • Data Processing and Smell Annotation: Java code was extracted from the Text‑Code subset of CodeXGLUE and parsed using Tree‑Sitter to annotate ten common code smell categories. This enabled the construction of parallel “smelly” and “clean” datasets. For each sample, we recorded token consumption and inference latency during both code repair and refactoring tasks using the DeepSeek‑R1 API.
  • Prompt Engineering and Optimization: To mitigate token overhead in CoT-based reasoning, we implemented and evaluated three prompt-level strategies: Context Awareness, Responsibility Tuning, and Cost Sensitive. Each strategy was applied by modifying the model’s input prompt, and for each experiment, we logged token usage, inference time, and code repair quality.
  • Automation and Analysis: All API interactions were automated via Python scripts, ensuring reproducibility and efficiency. Outputs—including token counts, timing, and repaired code—were collected and analyzed using custom scripts for data aggregation, statistical analysis, and visualization, enabling comprehensive comparison of different strategies.
IV Experimental Study: Triggering the nano surge
Figure 5: Nano Surge on Token Consumption
IV-A RQ1: Smelly Code vs. Clean Code